A rectangle has a dimension of 5 3\4 ft by 2 1\3 ft. What is the area and perimeter

Answer by rfer(16322)

5 3/4*2 1/3=
23/4*7/3=161/12=13 5/12 sq ft
----------------------------
2(5 3/4)+2(2 1/3)=
11 1/2+4 2/3=
138/12+56/12=194/12=16 1/6 ft